OpenStudy (anonymous):

-11x-10y=17 5x-7y=50 What do I do?

hero (hero):

Multiply both sides of the second equation by 2: Then you'll have the following system: -11x - 10y = 17 10x - 14y = 100 Afterwards combine them together: -x -24y = 117 Next isolate x: x = -24y - 117 Then substitute the expression for x into one of the original equations: 5(-24y - 117) -7y = 50 -120y - 585 - 7y = 50 Then continue solving for y -120y - 7y - 585 = 50 -(120 + 7)y = 585 + 50 -127y = 630 y = 635/-127 y = -5 From there, try solving for x

OpenStudy (anonymous):

does x = 17 ?

hero (hero):

No, \(x \ne 17\)

hero (hero):

\(-11x -10(-5) = 17\) \(-11x + 50 = 17\) \(-11x + 50 - 17 = 0\) \(-11 x+ 33 = 0\) \(33 = 11x\) \(\dfrac{33}{11} = x\)

OpenStudy (anonymous):

Oh ! I plugged in -5 to the second equation and I forgot two negatives make a positive so x will equal 3